# Sequential Quantum Measurements and the Instrumental Group Algebra

### v1

@@ I.2 The Commutative Analog
I.3 Summary of Contents and Main Results
  • I. Introduction
    §I.1. Prelude
    §I.2. The Commutative Analog
    §I.3. Summary of Contents
  • II. Quantum Measurement Theory: An Overview of Standard Knowledge
    §II.1. POVMs, Instruments, and Indirect Measurement
    §II.2. Total Operations (a.k.a. Channels) and Decoherence
    §II.3. Continuous-in-Time Decoherence/Dissipation
  • III. Continuous Measurements and Instrumental Groups
    §III.1. The Fundamental Problem of Continuous Measurement: Piling Kraus-Operators (Convolution 1 of 3)
    §III.2. Simultaneous Measurements of Non-Commuting Observables
    §III.3. Sequential Measurements: Kraus Operator Distributions and the Convolution (2 of 3)
  • IV. Kraus-Operator Densities, Ultraoperators, and the IG Algebra
    §IV.1. The Haar Measures: Kraus-Operator Densities and the Convolution (3 of 3)
    §IV.2. Translation Ultraoperators and their Invariant Derivatives
    §IV.3. The IG Algebra and Two Involutions: Gelfand-Kolmogorov and Cartan-Schmidt
    §IV.4. The Kolmogorov Equation for the KOD of a Continuous Measurement
  • V. Concluding Remarks
  • Appendix
    §A. System-Meter-Interaction Models
    §A.1. Jump Processes
    §A.2. Diffusive Processes
    §B. Invariant One-Forms, Haar Measures, and the Modular Function
    §C. The $\delta$-Function of a Group
Table 4: Table of Contents: A list of all the section titles for general reference.
Before explaining the main results of the paper, section II begins with a quick summary of all the basic tools that will be needed which are already familiar to quantum scientists — instruments, channels, and Lindbladians.
Section II.1 defines the generalized measurement, its Kraus operators, and the generalized Born rule. The positive-operator-valued measure (POVM) is introduced and the Luders rule is briefly mentioned. Certain operations are then considered, definining the quantum instrument and its subsequent total quantum operation. Then the all-important concept of an indirect measurement is discussed. The section finishes by addressing the position measurement.
Section II.2 then moves into the topic of decoherence channels and their usual superoperator properties, complete positivity and trace-preservation. The Hilbert-Schmidt adjoint is defined along with the Choi map and the subsequent superoperator adjoint dual to the Hilbert-Schmidt. Some care is then taken to ensure it is understood that the instrument behind the channel represents an observer and is not to be mistaken for an environment.
Section II.3 then introduces the most basic tools of continuous measurement. Lindblad operators and the Lindbladian superoperator are discussed. The one-parameter semigroup property is then explained as well as the subsequent Lindblad master equation. Although the stochastic Lindblad master equation should also be considered if we were interested in the evolution of the state during measurement, the subject is neglected because this paper is about instrument evolution.
Having reviewed the more-or-less standard theory that is required, we then move into newer concepts that have only recently emerged [3], but now from a convolution-based perspective.
With the dimension of time, measurements can be put together in a variety of ways — continuously, simultaneously, and sequentially — which can be considered to take place in an instrumental group (IG), the subject of section III:
Section III.1 explains the fundamental problem of continuous measurements, calculating the time-ordered product of a sequence of weak Kraus operators. It begins by immediately considering a sequence of two instruments as the convolution of their operation sets. The two weak instruments that generate the most common types of continuous Markov process are introduced, the jumpy Poisson process and the diffusive Wiener process. (Appendix A includes a detailed derivation of these two weak instruments from basic system-meter interactions.) The IG is then defined and the relation between the free product of groups and the convolution of operation sets is explained. Then the fundamental problem is stated more precisely and the topics of automation and chaos are briefly discussed.
Section III.2 explains the basic observation that the Kraus operators of weak measurements commute, even when the Lindblad operators that generate them do not. With this observation in mind, it then makes perfect sense to consider multiple weak measurements with noncommuting Lindblad operators as simultaneous. Such simultaneous measurements can be run for finite amounts of time, thus generating new types of instruments such as those generated by the general “DMNCOS” procedure. Critically, it is then explained that the finite-time instrument that is defined by such a procedure has its definition solely in the commutation relations of the Lindblad operators that generate their Kraus operators, therefore making the structure of the instrument universal, i.e. independent of Hilbert space. This emphasis is very important because it reinforces the basic fact that in physical examples, these commutation relations are based in physical observables that are more fundamental than the Hilbert spaces on which any of these Kraus operators act, a point that is echoes back to a point made by Kraus, referring to “ […] Bohr, who repeatedly stressed the importance of the classical nature of measuring instruments for the understanding of quantum mechanics.” [31, p. 2] A brief mention is then made to the IG for simultaneously measuring a canonical pair of phase-space quadratures and the IG for simultaneously measuring the three orthogonal angular momentum components and the reader is referred to previous work for further details.
Section III.3 then finishes with a fresh start, assuming an IG from the get-go and considering the instrument as a representation of the IG. This introduces the notion of a Kraus-operator distribution, a precursor to the Kraus-operator density that does not yet have the properties of a function. The convolution of two distributions over the IG is then defined and related to the operation-set convolution of two instruments in sequence. Finally, it is pointed out that this convolution of distributions already expands what it means for the total operations of a sequence of instruments to satisfy a group property.
Just like the real line, the IG is the domain of an entire suite of analytic tools — with invariant integrals, invariant derivatives, function spaces, and intertwined differential-algebraic equations — where functions such as the Kraus-operator density (KOD), the primary object of section IV, can be calculated, propogated, and varied:
Section IV.1 begins by introducing the all-important Haar integral measures. These measures allow one in turn to define the convolution of two functions and the density function of a distribution. The relationship between the convolution of functions and the convolution of distributions is explained. Finally, the definition of the Kraus-operator density (KOD) is fixed relative to the left-invariant measure, a choice that stems from the convention that time is considered to move from right to left when multiplying a sequence of Kraus operators.
Section IV.2 introduces the left- and right-translation ultraoperators. With them, convolution by a function can also be considered as an ultraoperator. The relations between these left- and right-convolution ultraoperators amongst themselves and each other are discussed. Then right- and left-invariant derivatives are defined. It is explained how these invariant derivatives generate the translation ultraoperators in the same analytic way that the usual derivative of a real line generates the Taylor expansion. Finally, ultraoperator commutators of the invariant derivatives are discussed in terms of the Lie algebra of the IG.
Section IV.3 discusses the function spaces that are defined by the IG and its left-invariant measure. Among the various norms one could give the IG-functional space, it is only the $\ell^{1}$ norm that is closed under convolution and the IG algebra (IGA) is finally defined as the Banach algebra of absolutely-integrable functions. The IG inner product is then introduced, the Kolmogorov adjoint is defined, and the Kolmogorov adjoints of the various ultraoperators are considered. Then the IGA is shown to have two involutions, named after Cartan and Gelfand, which respect the convolution product if the IG is unimodular. Finally, their relations to the Kolmogorov ultraoperator and Hilbert-Schmidt superoperator adjoints are discussed.
At very last, section IV.4 returns to the problem of continuous measurement and uses the ultraoperator formalism to rederive the KOD Kolmogorov equations in a style that is much like the reverse of section I.2. The various manifestations of the Markovian group property — either by composition, convolution, or addition — are summarized. The role of the instrument elements in intertwining the differential ultraoperators with the algebraic superoperators is then introduced. This fundamental intertwining relation is then used to formally derive the KOD Kolmogorov forward equation, showing along the way that the instrument elements also intertwine the Markov ultraoperator with the quantum channel superoperator. The Kolmogorov generator is then calculated for both diffusive and jump processes. Finally, it is shown that the instrument elements also intertwine both Kolmogorov generators with the Lindbladian dissipator, as one might expect, thus transforming the differential Kolmogorov equation into the algebraic Lindblad master equations, in noncommutative analogy with the Laplace transform.
The paper then finishes with some concluding remarks in section V.
For a table of contents, please refer to Table 4.
II Quantum Measurement Theory: An Overview of Standard Knowledge
Here is a summary of all the basic tools that will be needed which are already familiar to quantum scientists — instruments (II.1), channels (II.2), and Lindbladians (II.3).
II.1 Kraus operators, POVMs, and Instruments
$\sum_{x}\Omega_{x}^{\dagger}\Omega_{x}=1.$ (62)
$\mathcal{E}=\Big\{\Omega_{x}^{\dagger}\Omega_{x}\Big\}_{x}.$ (64)
$A\!\odot\!B^{\dagger}\,(\rho)\equiv A\rho B^{\dagger}.$ (66)
$\Omega_{x}=\sqrt{d\mu(x)}K_{x}.$ (70)
$\mathcal{E}=\{d\mu(x)K_{x}^{\dagger}K_{x}\}_{x}$ (71)
Without further context, the choice of how the Kraus-operator distribution is factored out of the instrument element is arbitrary or “ostensible” [44, 25]. In the context of continuous and sequential measurements, however, there is a distinct choice to make as will be explained in section III.3.
$\Omega_{x}=\langle x|\mathcal{U}_{\text{int}}|0\rangle,$ (73)
II.2 Total Operations (a.k.a. Channels) and Decoherence
the Choi involution
and the Choi-Jamiolkowski quasi-adjoint
@@ II.1 Kraus operators, POVMs, and Instruments
It is easy to then verify the following useful formulae
[…]
for which it is in turn easy to observe the properties
[…]
II.3 Continuous-in-Time Decoherence/Dissipation
$\mathrm{ad}[iH]\equiv-iH\odot 1+1\odot iH,$ (80)
[…]
where such an $\hat{L}$ is called a Lindblad operator [51]. General Lindbladians are the sum of a Hamiltonian superoperator and any number of dissipators. Dissipative channels are therefore generated by infinitesimal channels of the form
$\mathcal{Z}^{(\hat{L})}_{dt}\equiv e^{\mathcal{D}[\hat{L}]\kappa dt}$ (82)
$=e^{\mathcal{D}[\hat{L}]\kappa T}$ (84)
$\mathcal{Z}_{t+\Delta t}=\mathcal{Z}_{\Delta t}\circ\mathcal{Z}_{t},$ (85)
where the semigroup can be considered to refer to the set of channels $\{\mathcal{Z}_{t}\}_{t\geq 0}$, but it often can also refer implicitly to the semigroup of composite instrument elements as will be explained later (in section III.1, equation 93, and section III.3, equation 125). A state that evolves according to this channel $\rho_{t}=\mathcal{Z}_{t}(\rho_{0})$ satisfies a partial differential equation
III Continuous Measurement and Instrumental Groups
With the dimension of time, measurements can be put together in a variety of ways — continuously (III.1), simultaneously (III.2), and sequentially (III.3) — all of which can be considered to take place in an instrumental group (IG).
III.1 The Fundamental Problem of Continuous Measurement: Piling Kraus-Operators (Convolution 1 of 3)
@@
For the effective system-meter-interaction models that give 90 and 92, se Appendix A.
@@
where $\mathcal{I}^{*0}$ is the instrument with only one element, equal to the identity operation. As mentioned in section II, one should note that the convolution semigroup is quite a bit more elaborate than the one-parameter semigroup of total operations mentioned after equation 85. Indeed, the convolution semigroup is usually multi-dimensional and non-commutative. In practice, the convolution semigroup can often be contained within a group, what will be called the instrumental group (IG) and denoted $G_{\mathcal{I}}$. In section IV we will use the IG in a much more systematic way that begins with a more rigorous definition.
@@
up to some positive scalar factor $C\in\mathbb{R}_{>0}$ (which is added to the Kraus-operator distribution to be explained in section III.3.) This is more commonly known as the word problem [52], a topic that was very hot in the ’80s and ’90s, but mostly centered around digital ideas. Following [53], a group is said to be automatic if two products of group elements can be recognized as equal or distinct by a classical turing machine. The discussion in [53, 54, 55] was specifically interested in the much more complicated problem of finitely-generated groups, but the concept of an automatic group can just as well be applied to diffusively-generated groups. Here it will be left open exactly how to define what it means to “recognize” two chains as the same in general. In our cases of interest, the composite instrument element of every raw chain of events can be put into a certain canonical form that coordinatizes the IG. The ease in which these coordinatizations work thus suffices to demonstrate that such continuous IGs have the alleged form of automatic pattern recognition.
@@
III.2 Simultaneous Measurements of Non-Commuting Observables
@@
because the Kraus-operator distribution is a product of the individual Wiener measures for $dW$ and $dV$, and so the Wiener increments are independent. This means that even if the two Lindblad operators are noncommuting, their weak diffusive Kraus operators still commute. Thus two diffusive measurements can be done simultaneously, or (perhaps more descriptively) in tandem. Therefore given any finite number of potentially noncommuting Lindblad operators $\vec{L}=(\hat{L}_{1},\ldots,\hat{L}_{n})$ one can define the Diffusive Measurement of Non-Commuting Observables Simultaneously (DMNCOS) as the diffusive process generated by the weak instrument
@@
III.3 Sequential Measurements: Kraus-Operator Distributions and the Convolution (2 of 3)

### v2

@@ II A Commutative Analogy: The Real Line as an Instrument
III Prelude: Sequences of Measurement in Time and their Structure
III.1 There is a need to get past the idea of instantaneous measurement.
In quantum theory, observables and states are distinct objects, each a certain kind of operator over a Hilbert space. In particular circumstances, an observable can correspond to directly to a choice of Hermitian operator and the act of observation can be understood to be the random appearance of one of its eigenvalues with a probability that can be calculated by the Born rule. In this case, the appearance of the eigenvalue is also understood to signify that the state has projected onto the corresponding eigenspace. This idea of observation, in particular where the post-measurement state is projected, is called a von Neumann measurement or the Luders rule [22, 23, 24]. Useful as it is, the von Neumann-Luders rule fails to make sense of perhaps the most fundamental of observables, the position of a dynamic particle, because the eigendistributions of the position operator are not square-integrable and therefore not sensible state-preparations. (For a more detailed explanation, see the end of section A.1.)
III.2 With continuous measurements, what kinds of instruments are possible?
with
where $T$ is the overall time duration of the measurement, $\vec{L}\equiv(L_{1},L_{2},\ldots,L_{n})$ is a finite tuple of Lindblad operators (which generalize the Hermitian observable), $d\vec{W}_{[0,T)}\equiv\{d\vec{W}_{t}\}_{t=0}^{T-dt}$ is a sequence of Wiener increments (often called the measurement record) representing the raw data output by the instrument, $\mathcal{D}\mu[d\vec{W}_{[0,T)}]$ is the Wiener measure, and $\hat{X}$ denotes a representation of the abstract Lie algebra element $X$. (See Appendix B for a derivation of $\hat{\updelta}^{(L)}$.)
where
III.3 There is an analytic structure to applying instruments in sequence.
where $d_{\mathrm{L}}x$ is the left-invariant Haar measure, $\delta_{\text{IG}}(x)$ is the IG delta-function, and the dependence on any Lindblad operators has been dropped ($\gamma=\gamma^{\hat{L}}$). (See Appendix D for details on the delta-function.) What the KOD essentially does is count the number of measurement records that give the same Kraus operator $x$ at any given time $T$, therefore collecting the actions of a continuous measuring instrument.
$\mathcal{I}_{T}=\Big\{d_{L}x\,D_{T}(x)\mathcal{O}_{x}\Big\}_{x\in\text{IG}}$ (29)
III.4 There is a need to get past the state-centric approach to continuous measurement.
        Comparison of the KOD FPKE and the Lindblad master equation
V) KODs are in the IGA Density operators are in the dual of the von Neumann algebra
※(§V.4) Fokker-Planck-Kolmogorov equation Lindblad master equation
※(§V.4) Generator for a diffusive process
V.4) Generator for a jump process
V.4) Kolmogorov-Lindblad intertwining relation
IV.3) Instrument element irreps of the IG
or equivalently
is indeed a solution to equation 35. What then remains is the use of a few group-analytic tools which are explained in section V, particularly section V.4. The steps involve taking a derivative of the total operation with respect to time, applying the KOD FPKE, using the left-translation invariance of the measure ($d_{\mathrm{L}}(e^{hX}x)=d_{\mathrm{L}}x$) to move the right-invariant derivatives from the KOD to the superoperators $\mathcal{O}_{x}$, and finally using the Kraus-operator representation property ($K_{yx}=K_{y}K_{x}$) to observe the intertwining relations,
$\underleftarrow{L}[K_{x}]=\hat{L}K_{x}$ (38)
and therefore
III.5 The structure of sequential measurement is the IG convolution.
$\mathcal{Z}_{t}=e^{\mathcal{D}\kappa t}$ (40)
and has the well-known (one-parameter) group property
$\mathcal{Z}_{t+\Delta t}=\mathcal{Z}_{\Delta t}\circ\mathcal{Z}_{t}.$ (41)
where * denotes the IG convolution,
where $d_{\mathrm{R}}x$ is the right-invariant Haar measure of the IG. In terms of the right-invariant measure, this convolution can be recognized as the Chapman-Kolmogorov equation of a right-invariant Markov process. In terms of the left-invariant measure, this convolution allows the right-invariant Markov process to be converted into simple operator calculus, as will be explained in just a moment. Right-invariance is the equivalent (in IG language) to the property that the measuring process is nonadaptive, i.e. invariant to the Kraus operator that just preceeds the next step in time.
$\mathcal{Z}_{D}\equiv\int_{\text{IG}}d_{\mathrm{L}}y\,D(y)\mathcal{O}_{y},$ (44)
then for any second instrument with KOD $G(x)$
$\mathcal{Z}_{G}\circ\mathcal{Z}_{D}=\mathcal{Z}_{G*D}$ (45)
$\mathfrak{Z}^{\text{\Ankh}}_{t}=e^{\mathfrak{D}^{\text{\Ankh}}\kappa t}$ (48)
with the one-parameter group property
$\mathfrak{Z}^{\mathrm{L}}_{G}[D](x)\equiv G*D(x).$ (50)
$\mathfrak{L}_{g}[D](x)\equiv D(g^{\,\text{-}\hskip-1.0pt1}x).$ (52)
Summary of the Various Group Properties
From KODs to Ultraoperators From States to Superoperators
Invariant Chapman-Kolmogorov Equation
$D_{\Delta t}*D_{t}=D_{t+\Delta t}$$\Big|$
Instrumental Group Algebra
III.6 The home of the KOD is the IG algebra.
$G||_{1}\equiv\int_{\text{IG}}d_{\mathrm{L}}x\,|G(x)|,$ (55)
The superoperators $\mathcal{Z}_{f}$ defined in equation 44 and the hyperoperators $\mathfrak{Z}^{\mathrm{L}}_{f}$ defined in equation 53 are both representations of the IGA (by virtue of linearity and equations 45 and 51). However, between the two it is only the superoperators $\mathcal{Z}_{f}$ that compose a $\ddagger$}-representation and only the ultraoperators $\mathfrak{Z}^{\mathrm{L}}_{f}$ that compose an -representation. The representations of these involutions are the adjoints of the Hilbert-Schmidt and Kolmogorov inner products
@@ III.2 With continuous measurements, what kinds of instruments are possible?
There is apparently a canonical way to derive a C*-algebra in this case, the so-called enveloping C*-algebra [36, p. 47]. The exact meaning and correspondence of the enveloping $\text{C}^{\ddagger}$-algebra and the enveloping $\text{C}^{\text{\Ankh}}$-algebra require further investigations that are already beyond the purpose of the current paper.
IG convolutionV.1)
V.3) IG inner product: $(g,f)_{\text{IG}}\equiv\int_{\text{IG}}d_{\mathrm{L}}x\,\,\overline{g(x)}f(x)$ $(B|A)_{\mathcal{H}_{\lambda}}\equiv\operatorname{tr}_{\mathcal{H}_{\lambda}}B^{\dagger}A$ :Hilbert-Schmidt inner product
V.3) Kolmogorov adjoint: $\big(g,\mathfrak{Z}^{\text{\Ankh}}[f]\big)_{\text{IG}}\equiv\big(\mathfrak{Z}[g],f\big)_{\text{IG}}$ $\big(B\big|\mathcal{Z}^{\ddagger}(A)\big)_{\mathcal{H}_{\lambda}}\equiv\big(\mathcal{Z}(B)\big|A\big)_{\mathcal{H}_{\lambda}}$ :Hilbert-Schmidt adjoint
V.3) Gelfand involution: $f^{\text{\Ankh}}(x)\equiv\overline{f(x^{\,\text{-}\hskip-1.0pt1})}$ $f^{\ddagger}(x)\equiv\overline{f(x^{\dagger})}$ :Cartan involution (§V.3)※
V.2) left-translation ultraoperator: $\mathfrak{L}_{a}[f](x)\equiv f(a^{\,\text{-}\hskip-1.0pt1}x)$ $\mathcal{O}_{a}\equiv K_{a}\odot K_{a}^{\dagger}$ :CP superoperator (§IV.3)※
V.2) left-convolution ultraoperator: $\mathfrak{Z}^{\text{\Ankh}}_{f}\equiv\int_{\text{IG}}d_{\mathrm{L}}x\,\,f(x)\,\mathfrak{L}_{x}$ $\mathcal{Z}_{f}\equiv\int_{\text{IG}}d_{\mathrm{L}}x\,\,\overline{f(x)}\,\mathcal{O}_{x}$ :IGA superoperator (§V.3)※
IV Instrumental Groups: Moving from Continuous to Sequential Measurement
With the dimension of time, measurements can be put together in a variety of ways — continuously (IV.1), simultaneously (IV.2), and sequentially (IV.3) — all of which can be considered to take place in an instrumental group (IG).
IV.1 The Fundamental Problem of Continuous Measurement: Piling Kraus-Operators (Convolution 1 of 3)
@@
For the effective system-meter-interaction models that give 65 and 67, se Appendix B.
@@
where $\mathcal{I}^{*0}$ is the instrument with only one element, equal to the identity operation. As mentioned in section A, one should note that the convolution semigroup is quite a bit more elaborate than the one-parameter semigroup of total operations mentioned after equation 278. Indeed, the convolution semigroup is usually multi-dimensional and non-commutative. In practice, the convolution semigroup can often be contained within a group, what will be called the instrumental group (IG) and denoted $G_{\mathcal{I}}$. In section V we will use the IG in a much more systematic way that begins with a more rigorous definition.
@@
up to some positive scalar factor $C\in\mathbb{R}_{>0}$ (which is added to the Kraus-operator distribution to be explained in section IV.3.) This is more commonly known as the word problem [37], a topic that was very hot in the ’80s and ’90s, but mostly centered around digital ideas. Following [38], a group is said to be automatic if two products of group elements can be recognized as equal or distinct by a classical turing machine. The discussion in [38, 39, 40] was specifically interested in the much more complicated problem of finitely-generated groups, but the concept of an automatic group can just as well be applied to diffusively-generated groups. Here it will be left open exactly how to define what it means to “recognize” two chains as the same in general. In our cases of interest, the composite instrument element of every raw chain of events can be put into a certain canonical form that coordinatizes the IG. The ease in which these coordinatizations work thus suffices to demonstrate that such continuous IGs have the alleged form of automatic pattern recognition.
@@
IV.2 Simultaneous Measurements of Non-Commuting Observables
@@
because the Kraus-operator distribution is a product of the individual Wiener measures for $dW$ and $dV$, and so the Wiener increments are independent.
For readers who are not comfortable with either group commutators, Itô rules, or their combination, here is an alternative, more detailed explanation. Recall that an operator conjugated by an exponential can be calculated as an exponential series of nested commutators,
$e^{\hat{A}}\hat{B}e^{-\hat{A}}=\hat{B}+[\hat{A},\hat{B}]+\frac{1}{2}\big[\hat{A},[\hat{A},\hat{B}$$\big]$$\ldots$ (76)
and that operator conjugation commutes with exponentiation,
$ae^{\hat{B}}a^{\,\text{-}\hskip-1.0pt1}=e^{a\hat{B}a^{\,\text{-}\hskip-1.0pt1}}.$ (77)
In particular, for $\hat{A}=-\frac{1}{2}(\hat{L}^{\dagger}\hat{L}+\hat{L}^{2})\kappa dt+\hat{L}\sqrt{\kappa}dW$ and $\hat{B}=-\frac{1}{2}(\hat{M}^{\dagger}\hat{M}+\hat{M}^{2})\kappa dt+\hat{M}\sqrt{\kappa}dV$, where $dW$ and $dV$ are Wiener increments, we have
$[\hat{A},\hat{B}]=[\hat{L},\hat{M}]\kappa dWdV\hskip 50.0pt\text{and}\hskip 50.0pt\big[\hat{A},[\hat{A},\hat{B}]\big]=0$ (78)
where higher-order terms can be neglected by the Itô rules $dWdt=dVdt=0$, $dt^{2}=0$, and $dW^{2}=dt$. Finally, if $dW$ and $dV$ are independent, than even the last remaining term in the algebra commutator is zero by another Itô rule $dWdV=0$. Therefore even if $\hat{L},\hat{M}]\neq 0$, we still have $\hat{A},\hat{B}]=0$ and so the group elements represented by $K^{\hat{L}}(dW)$ and $K^{\hat{M}}(dV)$ also commute when $dW$ and $dV$ are independent. When $a$ and $b$ represent commuting group elements, it is not so natural to write an algebra commutator [b,a]=baab[b,a]=ba-ab because it requires defining subtraction. More natural is the group commutator $b^{\,\text{-}\hskip-1.0pt1}a^{\,\text{-}\hskip-1.0pt1}ba=1$, which is pictured as a sequence of 4 steps which return to the identity as in equation 75.
Equation 75 means that even if the two Lindblad operators are noncommuting, their weak diffusive Kraus operators still commute. Thus two diffusive measurements can be done simultaneously, or (perhaps more descriptively) in tandem. Therefore given any finite number of potentially noncommuting Lindblad operators $\vec{L}=(\hat{L}_{1},\ldots,\hat{L}_{n})$ one can define the Diffusive Measurement of Non-Commuting Observables Simultaneously (DMNCOS) as the diffusive process generated by the weak instrument
@@
IV.3 Sequential Measurements: Kraus-Operator Distributions and the Convolution (2 of 3)